NON-INTEGRAL TOROIDAL SURGERY ON HYPERBOLIC KNOTS IN S^{3}

CAMERON MCA. GORDON YING-QING WU XINGRU ZHANG

Abstract. We show that on any hyperbolic knot in S^{3} there is at most one nonintegral Dehn surgery which yields a manifold containing an incompressible torus.